Solve using the substitution method.
-x - 3y = -20
5x + 3y = 40

Answer by mananth(16946)

-1 x -3 y = -20 .............1
3 y = 40 -5 x ---------2
/ 3 =
y = 13.33 -1.67 x
Substitute y in (1)
-1x-3(13.33-1.67x)=-20
-1x-40+ 5x=-20
-1x+5x =-20+40
4x=20
/ 4
x= 5
Plug the value of x in (1)
-1 x -3 y = -20
-1 * 5 -3 y = -20
-5-3y=-20
-3y=-15
/-3
y=5